# Block-Attention for Efficient RAG

### v2

@@ 3.2 Efficiency
In the previous section, we already addressed our first concern: After fine-tuning, the Block-Attention model can achieve similar or even better performance than the self-attention model. In this section, we focus on our second concern: How much can the Block-Attention mechanism reduce the TTFT and the Flops of first token?
@@
To quantify the effects of the Block-Attention mechanism on the TTFT, we show in Table 2 the respective TTFTs and Flops of the Llama3-block and Llama3-vanilla when the KV states of all retrieved passages have been pre-computed and cached in memory. Obviously, the acceleration effect is gratifying: Once the length of the input sequence is 512 and the length of user input is 50, using Block-Attention can reduce TTFT by 48% and reduce Flops by 90.1%. As the total length increases, when it is equal to 32K, the acceleration effect reaches an astonishing 98.7%, and the consumption of Flops is even reduced by 99.8%. We may simply conclude the results as: with greater text comes greater necessity for Block-Attention.
Prompt Length 0 512 1K 2K 4K 8K 16K 32K
@@
Flops-vanilla 7.5e+11 7.6e+12 1.5e+13 3.0e+13 6.1e+13 1.2e+14 2.45e+14 4.9e+14
Flops-block 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11
@@ 3.3 Discussion
From the experimental results, we can easily figure out the effects of Block-Attention on existing RAG-based applications: Under the existing technical paradigm, developers need to deliberately balance the trade-off between accuracy and inference speed. Therefore, they have to limit the number of retrieved passages to a certain number, such as 3 to 10. With Block-Attention, the impact of the number of retrieved passages on inference speed will be greatly reduced, which will empower the developers to freely choose the number of retrieved passages that gives the optimal effect without any hesitation.
Furthermore, some dynamic RAG paradigms such as ReAct (Yao et al., 2023) have been proven to enhance the general task-solving ability of the model in most scenarios. However, compared to vanilla RAG, such paradigms need to iteratively compute the KV states of retrieved passages for the same user input, which results in a disastrous inference latency. With the help of Block-Attention, the additional latency consumption in this aspect can be reduced to a level that users can tolerate. Developers will be able to safely use such paradigms in online scenarios to optimize the reasoning ability of LLM.
@@ 4 Conclusion
Our experimental results across various RAG benchmarks demonstrated the profound impact of Block-Attention. We showed that Block-Attention can maintain the original reasoning accuracy of LLM while significantly reducing its TTFT and Flops in RAG scenarios. The effectiveness of Block-Attention becomes increasingly apparent as the number of passages increases or the frequency of retrievals increases. When considering the implementation of Block-Attention in your applications, bear in mind this guiding principle: With greater text comes greater necessity for Block-Attention.

### v3

@@ 3.6 Inference Efficiency of Block-Attention
In the previous section, we already addressed our first concern: After fine-tuning, the Block-Attention model can achieve similar or even better performance than the self-attention model. In this section, we focus on our second concern: How much can the Block-Attention mechanism reduce the TTFT and FLOPs-TFT?
@@
To quantify the effects of the Block-Attention mechanism on the efficiency, we show in Table 2 the respective TTFTs and FLOPs-TFT of the Llama3-block-ft and Llama3-vanilla-sft when the KV states of all retrieved passages have been pre-computed and cached in memory. Obviously, the acceleration effect is gratifying: Once the length of the input sequence is 512 and the length of user input is 50, using Block-Attention can reduce TTFT by 48% and reduce FLOPs-TFT by 90.1%. As the total length increases, the TTFT and FLOPs-TTF of the Block-Attention model maintain an essentially unchanged trend. When the length reaches 32K, the acceleration effect reaches an astonishing 98.7%, and the consumption of FLOPs-TFT is even reduced by 99.8%. We may simply conclude the results as: with greater text comes greater necessity for Block-Attention.
Prompt Length 50 512 1K 2K 4K 8K 16K 32K
@@
FLOPs-TFT-vanilla 7.5e+11 7.6e+12 1.5e+13 3.0e+13 6.1e+13 1.2e+14 2.45e+14 4.9e+14
FLOPs-TFT-block 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11 7.5e+11
@@ 3.7 Discussion
From the experimental results, we can figure out the effects of Block-Attention on existing RAG applications: Under the existing technical paradigm, developers need to deliberately balance the trade-off between accuracy and latency. Therefore, they have to limit the number of retrieved passages to a certain number, such as 3 to 10. With Block-Attention, the impact of the number of retrieved passages on inference speed will be greatly reduced, which will empower the developers to freely choose the number of retrieved passages that gives the optimal effect without any hesitation.
Furthermore, some dynamic RAG paradigms such as ReAct (Yao et al., 2023) have been proven to enhance the general task-solving ability of the LLM. However, compared to vanilla RAG, such paradigms need to iteratively compute the KV states of retrieved passages for single user input, which results in a disastrous inference latency. With the help of Block-Attention, the additional latency consumption in this aspect can be reduced to a level that users can tolerate. Developers will be able to safely use such paradigms in online scenarios to optimize the reasoning ability of LLM.
4 Related Works
4.1 Retrieval-Augmented Generation (RAG)
RAG is a widely used technique to improve generations of language models by using retrieved nearest-neighbor documents or passages as references, which typically involves two stages: retrieval and generation. Before generation, retrieval finds most similar passages with the user query or the context, by using BM25 or dense retrieval model (Lee et al., 2021; Lan et al., 2023; 2024). After collecting retrieved passages, there are numerous techniques to incorporate the knowledge during generation. Earlier works includes concatenation (Izacard & Grave, 2021) and cross-attention (Borgeaud et al., 2022). Some studies have also begun to attempt to directly use the retriever as text generator, that is, text generation is performed by selecting context-aware phrases from a collection of supporting documents.  (Lan et al., 2023; Cao et al., 2024)
Recently, LLMs becomes the most powerful paradigm for most NLP tasks, and simply concatenating all retrieved documents into the context of LLMs becomes the most simple and effective way for retrieval-augmented generation (RAG). For example, Self-RAG (Asai et al., 2023) leverage a critic model to decide which content in retrieved passages should be used during generation. As a specific application of RAG, tool learning is widely used to call external APIs to retrieve related passages from external database or tools to solve knowledge-intensive tasks (Schick et al., 2023).
4.2 Efficiency of RAG
The efficieny of LLM inference gradually decreases as the sequence length of the concatenation of retrieved passages rapidly increases, leading to a huge challenges for retrieval augmented generation of LLMs. To solve this problem, there are numerous novel architectures, model quantization and inference strategies are proposed. For example, the linear attention and Mixture of Expert (Fedus et al., 2022) architecture are widely used architectures for reducing the complexity of LLM inference (Shazeer, 2019; Zhang & Cai, 2022; Fedus et al., 2022). Speculative decoding leverages the parallel of language models to speed up the inference (Leviathan et al., 2023). However, most previous works are not designed for RAG scenarios. Recently, RAG Cache(Jin et al., 2024) caches the intermediate states of retrieved knowledge in GPUs, showing 4x TTFT.
Another stream of researches focus on individually and parallely process each documents, which is related to our work, such as FiD (Izacard & Grave, 2021), PCW (Ratner et al., 2023a) and Prompt Cache (Gim et al., 2024). However, FiD is widely used for encode-decoder architecture, and not compatible for the decoder-only architecture of LLMs, since they need to concatenate the hidden states for decoder during inference. PCW focuses on extending the context window rather than efficient inference. Prompt Cache enables the efficient inference for RAG applications by designing the Prompt Markup Language (PML), while their pre-computed KV cache of retrieved passages are bound to the specific IDs, leading to its limited applications. In other words, once the rank of a passage in the input sequence changes, its KV cache cannot be reused.
@@ 5 Conclusion
Our experimental results across various RAG benchmarks demonstrated the profound impact of Block-Attention. We showed that Block-Attention can maintain the original reasoning accuracy of LLM while significantly reducing its TTFT and FLOPs-TTF in RAG scenarios. The effectiveness of Block-Attention becomes increasingly apparent as the number of passages increases or the frequency of retrievals increases. When considering the implementation of Block-Attention in your applications, bear in mind this guiding principle: With greater text comes greater necessity for Block-Attention.